OpenStudy (anonymous):

Write an equation for the line passing throught the point (-8, 8) and having a slope of -1/2

OpenStudy (anonymous):

point slope form: y-8=-1/2(x+8)

OpenStudy (anonymous):

slope intercept form: y=-1/2x+4